Eduardo's school is selling tickets to a choral performance. On the first day of the ticket sales the school sold 9 senior citizen tickets and 2 student tickets for a total of $69. The school took in $128 on the second day by selling 14 senior citizen tickets and 10 student tickets. What is the price of each of one senior citizen ticket and one student ticket?
Answer by mananth(16949)

Senior citizens x unit price
student y numbers

9 x + 2 y = 69 .............1
14 x + 10 y = 128.00 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 1
-45 x -10 y = -345
14 x + 10 y = 128
Add the two equations
-31 x = -217.00
/ -31
x = 7
plug value of x in (1)
9 x + 2 y = 69
63 + 2 y = 69
2 y = 69
2 y = 6
y = 3
Senior citizens unit price 7
student unit price 3
m.ananth@hotmail.ca